how do you work this problem?
-5x-3y=-7
4x-y=-8

Answer by rfer(16322)

-5x-3y=-7
4x-y=-8
-----------------
y=4x+8
---------------
-5x-3(4x+8)=-7
-5x-12x-24=-7
-17x=17
x=-1
---------------
-5*-1-3y=-7
5-3y=-7
-3y=-12
y=4
-----------------
4*-1-4=-8